OpenStudy (anonymous):

How many three-digit numbers can be formed using the digits 0 and 7? Repeated digits are allowed.

OpenStudy (anonymous):

how many digit are there between 0 to 7?

OpenStudy (anonymous):

8

OpenStudy (amistre64):

making a tree diagram makes this thing huge

OpenStudy (anonymous):

but its only the digits 0 an7 so that doesnt matter how many are in between

OpenStudy (amistre64):

so we have to think of it logically; we have 8 choices; which have 8 choices, making 8*8 = 64 choices we have another 8 choices for each of those; 64*8 = 8^3 8*8*8 = 8^3 = number of options total

OpenStudy (anonymous):

Oh, it is just 0 and 7 so you have two choices for each digits 2 * 2 *2 =8

OpenStudy (amistre64):

think of a tree that has 8 branches, named 01234567 each branch has 8 more branches ... with 8 more branches extending from thos as well

OpenStudy (amistre64):

.... and the 0 i spose makes for complication lol

OpenStudy (anonymous):

its only the digits 0 and 7

OpenStudy (amistre64):

there are 000 to 099 that we can omit I spose?

OpenStudy (anonymous):

so its not as complicated as we think

OpenStudy (anonymous):

I believe my last response was right

OpenStudy (anonymous):

But we have two digits and 3 spaces ..

OpenStudy (anonymous):

why would it be 2 * 2* 2?

OpenStudy (amistre64):

lol ... this belongs in the writing section ;) maybe

OpenStudy (anonymous):

two choices for first digit 0,7 two choices for second digits 0,7 two choices for third digits 0,7 2*2*2=8

OpenStudy (amistre64):

000 007 070 700 077 707 770 777 is that all of them?

OpenStudy (anonymous):

yeah got it

OpenStudy (amistre64):

aint that 3C2 as well?

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

Not combination or permutation because numbers are allowed to repeat

OpenStudy (anonymous):

yeah